OpenStudy (anonymous):

PLEASE HELP!!!!!! An equation is shown below: km^2 - c = n Solve the equation for m.

OpenStudy (anonymous):

@kagıtucak

OpenStudy (sidsiddhartha):

\[km^2-c=n\] right?

OpenStudy (anonymous):

yes. sorry, i'll fix that

OpenStudy (anonymous):

Add C. then you have km2=n+c. Divide by K. then have m2=n+c/k, THEN Take the Sqrt. then get\[m=\sqrt{n+c/k}\]

OpenStudy (sidsiddhartha):

first just take c on the other side \[km^2=n+c\] ok?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

wait, so is this the answer @MarrBear ?

OpenStudy (anonymous):

YES IT IS! xD

OpenStudy (sidsiddhartha):

now divide both sides by k \[\frac{ km^2 }{ k }=\frac{ n+c }{ k }\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Anytime c:

OpenStudy (sidsiddhartha):

u'll have \[m^2=(n+c)/k \rightarrow m=\sqrt{(n+c)/k}\]

OpenStudy (anonymous):

ok :) thanks you so much 2 @sidsiddhartha :) wish i could give u a medal 2 :) sorry :'(

OpenStudy (sidsiddhartha):

no problem medals not everything ^_^

OpenStudy (anonymous):

:)